OpenStudy (anonymous):

Find the equation of a parabola passing through (-4, -6) whose vertex is at (-2,-4) and whose axis is parallel to the x-axis. the answer is (y+2)^2 = -2(x+2).. how to get the P given the points, vertex? Help me please. Show it to me step by step. Thank you.

OpenStudy (anonymous):

r u sure its the right question and what do you mean by P

OpenStudy (anonymous):

The answer is (y+4)^2 = -2(x+2).. It wasn't (y+2)^2

OpenStudy (anonymous):

P is the distance between the vertex and the focus, and also the same (that is, equal) distance between the vertex and the directrix..

OpenStudy (jdoe0001):

hmmm how do you know "the answer is (y+2)^2 = -2(x+2)"

OpenStudy (anonymous):

it was written in the book and I don't know how to get the answer..

OpenStudy (jdoe0001):

something tells me you you've got the wrong answer :/

OpenStudy (jdoe0001):

anyhow \(\bf (y-k)^2=4p(x-h)\) (h, k) = vertex point p = distance of the focus from the vertex

OpenStudy (anonymous):

yes thats the formula..

OpenStudy (anonymous):

the vertex's given..

OpenStudy (jdoe0001):

for an equation of say => (y+2)^2 = -2(x+2) then the 4P = -2 solve for P to get the distance from the vertex though notice that the (h, k) in that equation is (-2, -2)

OpenStudy (anonymous):

The answer is (y+4)^2 = -2(x+2)

OpenStudy (anonymous):

its not (y+2)^2.. sorry

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

How do I get p given the vertex and the points??